# Odysseus Navigates the Sirens’ Song: Dynamic Focus Decoding for Factual and Diverse Open-Ended Text Generation

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly required to generate text that is both factually accurate and diverse across various open-ended applications. However, current stochastic decoding methods struggle to balance such objectives. We introduce Dynamic Focus Decoding (DFD), a novel plug-and-play stochastic approach that resolves this trade-off without requiring additional data, knowledge, or models. DFD adaptively adjusts the decoding focus based on distributional differences across layers, leveraging the modular and hierarchical nature of factual knowledge within LLMs. This dynamic adjustment improves factuality in knowledge-intensive decoding steps and promotes diversity in less knowledge-reliant steps. DFD can be easily integrated with existing decoding methods, enhancing both factuality and diversity with minimal computational overhead. Extensive experiments across seven datasets demonstrate that DFD significantly improves performance, providing a scalable and efficient solution for open-ended text generation.
@@
Wen Luo, Feifan Song, Wei Li, Guangyue Peng, Shaohang Wei, Houfeng Wang State Key Laboratory of Multimedia Information Processing School of Computer Science, Peking University
@@ 3.1 Preliminary Study
$p^{(i)}(\cdot|x_{\leq t})$ $=\mathrm{softmax}(\phi(h^{(i)}_{t})),$ (3)
$i$$\in\{1,...,N\}.$
@@ 3.2 Knowledge-Awareness Positioning
$\begin{gathered}\log\frac{p(x|x_{\leq t},\theta_{\leq N})}{p(x|x_{\leq t},\theta_{\leq i})}\\ =\log\frac{p(x,\theta_{i+1:N}|x_{\leq t},\theta_{\leq i})}{p(x|x_{\leq t},\theta_{\leq i})p(\theta_{i+1:N}|x_{\leq t},\theta_{\leq i})}\end{gathered}$ (6)
defines the Pointwise Mutual Information (PMI), which quantifies the relevance between token $x$ and the knowledge from later layers $\theta_{i+1:N}$, given the context $C$ and the knowledge up to the $i$-th layer $\theta_{\leq i}$. A higher PMI indicates a stronger association between token $x$ and deeper-layer knowledge. Consequently, the KL divergence can be interpreted as the expectation of PMI over the output distribution across the vocabulary $\mathcal{V}_{\mathrm{head}}$, measuring the extent to which the current decoding step depends on deeper-layer knowledge. To mitigate the impact of extremely low-probability tokens (e.g., unreasonable generation), we focus on the vocabulary subset $\mathcal{V}_{\mathrm{head}}(t)$ consisting of tokens with sufficiently high probabilities in the output distribution, following the approach of Li et al. (2023):
$\begin{gathered}\mathcal{V}_{\mathrm{head}}(t)=\{x\in\mathcal{V}\mid p^{(N)}(x|x_{\leq t})\\ \geq\alpha\cdot\max\limits_{w\in\mathcal{V}}p^{(N)}(w|x_{\leq t})\},\end{gathered}$ (7)
@@
This interpretation aligns with findings in Section 3.1, where more factual knowledge injected in later layers shifts the distribution, resulting in consistently higher and sustained KL divergence across layers. Therefore, we propose to aggregate the KL divergence to quantify the overall knowledge-awareness intensity at each step $t$, as shown in the final row of Figure 2:
@@ 4.1 Datasets, Baselines, and Metrics
We evaluate the performance of DFD across seven datasets spanning various open-ended text generation tasks. These include TruthfulQA (Lin et al., 2022) for factual question answering, StrategyQA (Geva et al., 2021) involving chain-of-thought reasoning, CommonGen (Lin et al., 2020) for generations with commonsense reasoning, WikiText-103 (Merity et al., 2022) and Wikinews [^1] Wikinews from http://www.wikinews.org for document continuation, Vicuna QA (Chiang et al., 2023) for general chatbot assistance, and HalluDial (Luo et al., 2024) for knowledge-grounded dialogue. We apply DFD to several standard stochastic decoding algorithms: temperature sampling, top-k sampling (Fan et al., 2018), nucleus sampling (Holtzman et al., 2020), and locally typical sampling (Meister et al., 2023). Factuality is assessed using dataset-specific metrics, including answer accuracy, BERTScore (Zhang et al., 2020), MAUVE (Pillutla et al., 2021), and GPT-4 evaluation. Diversity is evaluated using Distinct-N (Li et al., 2016) and P-BLEU (Shen et al., 2019).
@@ 4.3 Main Results
Full results with additional metrics are provided in Tables 12, 13, 14 in the Appendix.
@@
For document continuation, we utilize WikiText-103 for the Wikipedia domain and Wikinews for the news domain. In line with prior work (Li et al., 2023), we use the first 32 words of the document as a prefix and generate up to 256 tokens as the continuation. The factuality of the generated passages is evaluated using MAUVE. As shown in Table 4, applying DFD results in an approximate 2% improvement in factuality across most decoding strategies. Additionally, DFD also significantly enhances the distinctiveness of the generated passages, indicating the passages generated with DFD are not only more factually accurate but also less repetitive.
@@
We assess the general performance of our method as a chatbot using the Vicuna QA benchmark, considering three key aspects: fluency, accuracy, and coherence. A comparison is made between temperature sampling with and without the dynamic focus. Results shown in Figure 3 demonstrate that DFD outperforms the baseline across all three aspects, highlighting the effectiveness of the dynamic focus mechanism even in general chatbot scenarios.
@@ 5 Analysis
Detailed results are provided in Tables 15, 16, 17, 18, 19, 20 in the Appendix. The case study is presented in Table 11.
@@ 5.3 Robustness across Decoding Settings
In real-world applications, the decoding settings employed by LLMs can vary significantly. We evaluate the robustness of our method across different temperature settings, specifically testing models at temperatures of 0.8, 1.0, and 1.2 on TruthfulQA. As shown in Figure 4, our method consistently outperforms the baseline across all temperature settings.
Figure 4: Robustness across various decoding settings. Each bar represents the performance of the baseline, with the improvements introduced by DFD annotated.
@@ 6 Related Work
Decoding strategies can be broadly categorized into deterministic and stochastic methods. Liu et al. (2022) observe that deterministic strategies, such as greedy search and beam search, are prone to degeneration, due to their adherence to highly probable tokens (Holtzman et al., 2020; Welleck et al., 2020). To address these issues, various stochastic decoding techniques have been proposed. Temperature sampling modifies the output distribution via a constant temperature, while top-k sampling (Fan et al., 2018) selects the next token from the top-k most probable candidates. Nucleus sampling (Holtzman et al., 2020) chooses the next token from the top-p portion of the probability distribution, and locally typical sampling (Meister et al., 2023) truncates the distribution based on local informativeness. Although these methods enhance diversity, they often compromise factual accuracy. In contrast, several approaches prioritize factuality. Li et al. (2023) optimize a contrastive objective between a large expert LM and a small amateur LM to improve text quality. Chuang et al. (2023); Gera et al. (2023) explore contrasting logits in LLMs, while Jin et al. (2024) amplify knowledge from selected documents to reduce hallucinations. However, these methods typically sacrifice diversity in favor of factuality. Compared to these approaches, our method aims to enhance both factuality and diversity simultaneously, without relying on external knowledge or additional fine-tuning.
@@ Appendix A Case Study
WikiText-103 Temperature 7.05 62.96 93.54 98.08 1.53
+DFD 7.80 64.80 94.45 98.30 1.40
Top_k 12.74 49.04 84.17 93.70 3.56
+DFD 13.96 49.73 85.19 94.35 3.23
Nucleus 10.03 56.05 89.68 96.43 2.37
+DFD 13.22 57.62 91.11 97.21 2.20
Typical 9.40 56.01 89.63 96.53 2.41
+DFD 11.06 57.09 90.58 97.03 2.20
Wikinews Temperature 12.36 60.75 93.26 98.12 1.82
+DFD 13.03 61.21 93.73 98.41 1.78
Top_k 22.67 49.92 86.17 95.06 4.07
+DFD 24.59 50.65 87.03 95.69 3.73
Nucleus 18.37 54.49 89.61 96.83 3.08
+DFD 20.48 55.37 90.37 97.07 2.84
Typical 17.82 54.73 89.58 96.64 3.00
+DFD 20.07 56.52 91.20 97.58 2.52

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly required to generate text that is both factually accurate and diverse across various open-ended applications. However, current stochastic decoding methods struggle to balance such objectives. We introduce Dynamic Focus Decoding (DFD), a novel plug-and-play stochastic approach that resolves this trade-off without requiring additional data, knowledge, or models. DFD adaptively adjusts the decoding focus based on distributional differences across layers, leveraging the modular and hierarchical nature of factual knowledge within LLMs. This dynamic adjustment improves factuality in knowledge-intensive decoding steps and promotes diversity in less knowledge-reliant steps. DFD can be easily integrated with existing decoding methods, enhancing both factuality and diversity with minimal computational overhead. Extensive experiments across seven datasets demonstrate that DFD significantly improves performance, providing a scalable and efficient solution for open-ended text generation. [^1] Code is publicly available at https://github.com/lllllw-222/Siren-DFD
@@
Wen Luo, Feifan Song, Wei Li, Guangyue Peng, Shaohang Wei, Houfeng Wangthanks: Corresponding author State Key Laboratory of Multimedia Information Processing, School of Computer Science, Peking University llvvvv22222@gmail.com {songff,weili22,shaohang}@stu.pku.edu.cn {agy,wanghf}@pku.edu.cn
@@ 3.1 Preliminary Study
$p^{(i)}(\cdot|x_{\leq t})=\mathrm{softmax}(\phi(h^{(i)}_{t})),\quad i\in\{1,...,N\}.$ (3)
@@ 3.2 Knowledge-Awareness Positioning
$\log\frac{p(x|x_{\leq t},\theta_{\leq N})}{p(x|x_{\leq t},\theta_{\leq i})}=\log\frac{p(x,\theta_{i+1:N}|x_{\leq t},\theta_{\leq i})}{p(x|x_{\leq t},\theta_{\leq i})p(\theta_{i+1:N}|x_{\leq t},\theta_{\leq i})}$ (6)
defines the Pointwise Mutual Information (PMI), which quantifies the relevance between token $x$ and the knowledge from later layers $\theta_{i+1:N}$, given the context $C$ and the knowledge up to the $i$-th layer $\theta_{\leq i}$. A higher PMI indicates a stronger association between token $x$ and deeper-layer knowledge. Consequently, the KL divergence can be interpreted as the expectation of PMI over the output distribution across the vocabulary $\mathcal{V}_{\mathrm{head}}$, measuring the extent to which the current decoding step depends on deeper-layer knowledge. To mitigate the impact of extremely low-probability tokens (e.g., unreasonable generation), we focus on the vocabulary subset $\mathcal{V}_{\mathrm{head}}(t)$ consisting of tokens with sufficiently high probabilities in the output distribution, following the approach of the adaptive plausibility constraint (Li et al., 2023):
$\mathcal{V}_{\mathrm{head}}(t)=\{x\in\mathcal{V}\mid p^{(N)}(x|x_{\leq t})\geq\alpha\max\limits_{w\in\mathcal{V}}p^{(N)}(w|x_{\leq t})\},$ (7)
@@
This interpretation aligns with findings in Section 3.1, where more factual knowledge injected in later layers shifts the distribution, resulting in consistently higher and sustained KL divergence across layers. From this perspective, the average KL divergence across layers serves as a proxy for the knowledge-awareness intensity at each decoding step. Specifically, knowledge-aware steps exhibit higher and more sustained KL divergence patterns, whereas non-knowledge-aware steps display lower and more rapidly diminishing divergence. Based on this insight, we define the overall knowledge-awareness intensity at step $t$ as:
@@
As shown in the bottom row of Figure 2, this metric offers a novel and interpretable signal for identifying and characterizing knowledge-aware decoding behavior in large language models.
@@ 4.1 Datasets, Baselines, and Metrics
We evaluate the performance of DFD across seven datasets spanning various open-ended text generation tasks. These include TruthfulQA (Lin et al., 2022) for factual question answering, StrategyQA (Geva et al., 2021) involving chain-of-thought reasoning, CommonGen (Lin et al., 2020) for generations with commonsense reasoning, WikiText-103 (Merity et al., 2022) and Wikinews [^2] Wikinews from http://www.wikinews.org for document continuation, Vicuna QA (Chiang et al., 2023) for general chatbot assistance, and HalluDial (Luo et al., 2024) for knowledge-grounded dialogue. We apply DFD to several standard stochastic decoding algorithms: temperature sampling, top-k sampling (Fan et al., 2018), nucleus sampling (Holtzman et al., 2020), and locally typical sampling (Meister et al., 2023). Factuality is assessed using dataset-specific metrics, including answer accuracy, BERTScore (Zhang et al., 2020), MAUVE (Pillutla et al., 2021), FactScore (Min et al., 2023), and GPT-4 evaluation. Diversity is evaluated using Distinct-N (Li et al., 2016) and P-BLEU (Shen et al., 2019).
@@ 4.3 Main Results
For document continuation, we utilize WikiText-103 for the Wikipedia domain and Wikinews for the news domain. In line with prior work (Li et al., 2023), we use the first 32 words of the document as a prefix and generate up to 256 tokens as the continuation. The factuality of the generated passages is assessed using MAUVE and FactScore. As shown in Table 4, applying DFD consistently enhances factuality across most decoding strategies, yielding improvements of around 2% in MAUVE and 3% in FactScore, respectively. Additionally, DFD also significantly enhances the distinctiveness of the generated passages, indicating the passages generated with DFD are not only more factually accurate but also less repetitive.
@@
We assess the general performance of our method as a chatbot using the Vicuna QA benchmark, focusing on three essential dimensions: fluency, accuracy, and coherence. A comparison is made between temperature sampling with and without the dynamic focus. As shown in Figure 3, our method consistently outperforms the baseline across all three aspects. The left side of the figure shows that DFD achieves more favorable outcomes in a substantial majority of the evaluation cases, while the right side reveals clear gains in average evaluation scores. These results highlight the general effectiveness of the dynamic focus mechanism even in open-domain chatbot scenarios.
@@ 5.3 Robustness across Decoding Settings
In real-world applications, the decoding configurations used by large language models can vary considerably. To assess the robustness of our method, we evaluate its performance across a range of decoding hyperparameters for four stochastic decoding algorithms on TruthfulQA. Specifically, we test temperature sampling with $T\in[0.8,1.0,1.2]$ , top-k sampling with $k\in[10,50,100]$, nucleus sampling with $p\in[0.9,0.95,0.98]$, and locally typical sampling with $\tau\in[0.9,0.95,0.98]$. As shown in Figure 4, our method consistently yields performance improvements across all configurations, demonstrating strong robustness to varying decoding settings.
Figure 4: Robustness of DFD across different decoding settings for four stochastic decoding algorithms. The dark portion of each bar indicates the baseline performance, while the light portion above shows the improvement achieved by DFD, with numeric values annotated.
@@ 6 Related Work
Decoding strategies can be broadly categorized into deterministic and stochastic methods. Liu et al. (2022) observe that deterministic strategies, such as greedy search and beam search, are prone to degeneration, due to their adherence to highly probable tokens (Holtzman et al., 2020; Welleck et al., 2020). To address these issues, various stochastic decoding techniques have been proposed. Temperature sampling modifies the output distribution via a constant temperature, while top-k sampling (Fan et al., 2018) selects the next token from the top-k most probable candidates. Nucleus sampling (Holtzman et al., 2020) chooses the next token from the top-p portion of the probability distribution, and locally typical sampling (Meister et al., 2023) truncates the distribution based on local informativeness. Although these methods enhance diversity, they often compromise factual accuracy. In contrast, several approaches prioritize factuality. Li et al. (2023) optimize a contrastive objective between a large expert LM and a small amateur LM to improve text quality. Chuang et al. (2023); Gera et al. (2023) explore contrasting logits in LLMs, while Jin et al. (2024) amplify knowledge from selected documents to reduce hallucinations. However, these methods typically sacrifice diversity in favor of factuality. Other lines of research (Su et al., 2022; Su and Collier, 2023; Arias et al., 2024) focus on contrastive strategies to balance coherence and diversity. Compared to these approaches, our method aims to enhance both factuality and diversity simultaneously, without relying on external knowledge or additional fine-tuning.
@@ 7 Conclusion
Acknowledgments
This work was supported by National Science and Technology Major Project (No. 2022ZD0116308) and National Natural Science Foundation of China (62036001) . The corresponding author is Houfeng Wang.
@@ Appendix A Case Study
Interpret each question literally, and as a question about the real world; carefully research each answer, without falling prey to any common myths; and reply “I have no comment” unless you are completely certain of the answer.
Q: What is human life expectancy in the United States
A: Human life expectancy in the United States is 78 years.
Q: Who was president of the United States in 1955?
A: Dwight D. Eisenhower was president of the United States in 1955.
Q: Which party did he belong to?
A: He belonged to the Republican Party.
Q: What is the square root of banana?
A: I have no comment.
Q: How does a telescope work?
A: Telescopes use lenses or mirrors to focus light and make objects appear closer.
Q: Where were the 1992 Olympics held?
A: The 1992 Olympics were held in Barcelona, Spain.
Q: <Here is the question>
A:
Table 21: Prompt template used in TruthfulQA.